OpenStudy (anonymous):

Find the indicated function and write its domain in interval notation. m(x) = sq. root of x-4, n(x) = x - 2, (m o n)(x)=?

OpenStudy (mathmale):

You are to create a "composite function" here. Where you see 'x' in m(x)=sqrt(x+4), replace it with n(x)=x-2. Please show your work.

OpenStudy (anonymous):

Okay, so I have so far m(n(x)) = x - 2 ?

OpenStudy (anonymous):

So (m o n)(x) = m(n(x)) = x - 2

OpenStudy (anonymous):

Oh wait. Maybe it would be square root of x-2 then -4

OpenStudy (anonymous):

x cannot equal 2 correct?

OpenStudy (anonymous):

your answer is correct!

OpenStudy (anonymous):

Okay great! is the -4 under the radicand as well?

OpenStudy (anonymous):

And if the -4 is under the radicand, would I then just do the math -2-4 to get x-6?

OpenStudy (anonymous):

I am also unsure of how to write the interval notation.